Here is the following corporation data:

Expected sales = $2,500,000,

Expected variable costs = $1,300,000

Expected fixed costs = $240,000.  

What is the breakeven in sales dollars?
Answer #1

Answer: Breakeven in sales dollars= $461,538.46

Explanation:

Breakeven point is the level at which total sales revenue=total expenses(Fixed costs+variable costs).

Breakeven point in $=Fixed Costs/Gross profit margin

Gross profit margin=variable costs/sales=1,300,000/2,500,000=0.52

putting values in formula;

=240,000/0.52=$461,538.46

If helpful, Thumbs UP please:)